Prove that 3n+4n+5n is divisible by 12 whenever n is an odd positive integer.


1
Expert's answer
2021-12-06T16:26:38-0500

As i understand it should be "3^n+4^n+5^n"

To prove that number is divisible by 12 we will prove that it is divisible by both 3 and 4


Divisible by 3:

"3^n" is obviously divisible by 3 for every odd integer n, then the point is to prove that "4^n+5^n" is divisible by 3

According to abbreviated multiplication formula for odd degrees:

"4^n+5^n=(4+5)(4^{n-1}-4^{n-2}*5+4^{n-3}*5^2-...-4*5^{n-2}+5^{n-1})=9*(4^{n-1}-4^{n-2}*5+4^{n-3}*5^2-...-4*5^{n-2}+5^{n-1})\u22ee3"

Divisible by 4:

"4^n" is obviously divisible by 4 for every odd integer n, then the point is to prove that "3^n+5^n" is divisible by 4

According to abbreviated multiplication formula for odd degrees:

"3^n+5^n=(3+5)(3^{n-1}-3^{n-2}*5+3^{n-3}*5^2-...-3*5^{n-2}+5^{n-1})=8*(3^{n-1}-3^{n-2}*5+3^{n-3}*5^2-...-3*5^{n-2}+5^{n-1})\u22ee4"


The statement has been proven